In recent years, Crane Company has purchased three machines. Because of frequent employee turnover in the accounting department, a different accountant was in charge of selecting the depreciation method for each machine, and various methods have been used. Information concerning the machines is summarized in the table below.

Machine

Acquired

Cost

Salvage
Value

Useful Life
(in years)

Depreciation
Method

1

Jan. 1, 2015 $126,000 $16,000 10 Straight-line

2

July 1, 2016 79,000 11,200 5 Declining-balance

3

Nov. 1, 2016 71,900 7,900 6 Units-of-activity


For the declining-balance method, Crane Company uses the double-declining rate. For the units-of-activity method, total machine hours are expected to be 32,000. Actual hours of use in the first 3 years were: 2016, 810; 2017, 6,400; and 2018, 7,900.

(a)

Correct answer iconYour answer is correct.

Compute the amount of accumulated depreciation on each machine at December 31, 2018.

MACHINE 1

MACHINE 2

MACHINE 3

Accumulated Depreciation at December 31

$44000

$56248

$30220

(b)

If machine 2 was purchased on April 1 instead of July 1, what would be the depreciation expense for this machine in 2016? In 2017?

2016

2017

Depreciation Expense

$enter a dollar amount

$enter a dollar amount
Answer #1

a) Depreciation expense per unit = (71900-7900/32000) = 2

b) Accumulated depreciation

Machine 1 Machine 2 Machine 3
Accumulated depreciation at december 31 (126000-16000/10*4) = 44000 (79000*40%+79000*60%*40%+79000*60%*60%*40%*6/12) = 56248 (810+6400+7900)*2 = 30220

c)

2016 2017
Depreciation expense 79000*40%*9/12 = 23700 79000*40%*3/12+79000*60%*40%*9/12 = 22120